OpenStudy (anonymous):

Solve by substitution 14x-35=7y -25-6x=5y

OpenStudy (anonymous):

[1] 14x = 7y + 35 [1] x = y/2 + 5/2 // Plug this in for variable x in equation [2] [2] -6•(y /2+5/2) - 5y = 25 [2] - 8y = 40 // Solve equation [2] for the variable y [2] 8y = - 40 [2] y = - 5 // By now we know this much : x = y/2+5/2 y = -5 // Use the y value to solve for x x = (/2)(-5)+5/2 = 0 Solution : {x,y} = {0,-5}

OpenStudy (anonymous):

{x,y} = {0,-5}

OpenStudy (anonymous):

thank you so much. do you think you can help me with another?